# RAMBO: Reliability Analysis for Mamba through Bit-flip attack Optimization

### v1

@@ Abstract.
Among such threats, hardware-induced bit-flip attacks (BFAs) pose a particularly severe risk by corrupting model parameters through memory faults, thereby undermining model accuracy and functional integrity. To investigate this vulnerability, we introduce RAMBO, the first BFA framework specifically designed to target Mamba-based architectures. Through experiments on the Mamba-1.4b model with LAMBADA benchmark—a cloze-style word- prediction task—we demonstrate that flipping merely a single critical bit can catastrophically reduce accuracy from 74.64% to 0% and increase perplexity from 18.94 to 3.75 × 106. These results demonstrate the pronounced fragility of SSMs to adversarial perturbations. The framework is open-sourced at https://anonymous.4open.science/r/RAMBO-DA22.
@@ 1. Introduction
State-space architectures such as Mamba implement fundamentally different information-propagation and parameterization mechanisms, trade off recurrence and selectivity for attention mechanisms (gu2023Mamba). Due to these architectural distinctions, it is not appropriate to assume that attack strategies and defenses developed for transformers transfer directly to SSMs. Consequently, the absence of a systematic study of BFAs on SSMs constitutes a substantive gap in the current AI hardware robustness and security literature. To address this gap, we propose the framework “RAMBO” — a first of its kind SSM-aware BFA pipeline. The primary contribution of the paper are as follows:
  • We identify and formalize a previously unexplored vulnerability of state-space models (e.g. Mamba) to bit-flip attacks, and propose RAMBO, a first of its kind SSM-aware attack approach, bridging the gap between BFA research and structured sequence models.
  • RAMBO leverages the structural properties of Mamba-style SSM layers to prioritize critical parameter regions, adapts gradient-estimation and search heuristics, and uses the resulting perturbation effects to identify minimal bit-flip sets that maximally disrupt model behavior.
  • RAMBO uncovers a significant vulnerability of Mamba models. A mere one bit-flip (7.14 × 10-10% of all bits) in Mamba-1.4b, can reduce the LAMBADA word prediction accuracy from 74.64% to 0%, while increasing Wikitext perplexity from 18.94 to 3.75 × 106.
@@ 3. Threat Model
In this work, we consider both white-box and gray-box threat models. The white-box setting applies because the targeted models are open-sourced, granting full access to their architecture and parameters. The gray-box setting reflects scenarios in which the adversary has only partial visibility into model internals. Within this setting, we focus on untargeted attacks, which aim to induce graceless degradation in overall model performance rather than targeting specific outputs. Such attacks are particularly insidious, as they compromise accuracy across a broad range of inputs while avoiding distinct failure patterns, making them more challenging to detect and defend against than targeted attacks (codematch; qian2023survey; nazari2024forget). Therefore, RAMBO strategically employs an untargeted attack to maximize disruption and degrade model performance. This attack also fundamentally differs from denial of service attacks, which aim to overwhelm system resources to render services unavailable.
4. Proposed RAMBO Methodology
@@ 4.1. Theoretical Vulnerability Analysis
Here, we present parameter-ranking procedure to identify plausible targets for the RAMBO attack. We rank the SSM parameters’ based on their importance as follows:
@@ 5.1. Experimental Setup
We evaluated RAMBO on a diverse set of models, including Mamba and Mamba2, ranging from 370 million and 2.8 billion parameters (Mamba; Mamba2). We further assess RAMBO on quantized models such as Quamba-1.4b-w8a8 (8-bit weights and 8-bit activations), Quamba-1.4b-w4a16, Quamba-2.8B-w8a8, and Quamba-2.8b-w4a16 to evaluate its effectiveness under low-precision settings. Model performance was assessed using standard benchmarks, such as the tasks from the Language Model Evaluation Harness (eval-harness), including Arc-Easy, HellaSwag, PIQA and Winograde, which probe reasoning and generalization across a variety of domains. Additionally, we tested on the LAMBADA dataset (paperno2016lambada), a word-prediction/ cloze-style natural language understanding task. We extended our evaluation to include Vision-Mamba models (hatamizadeh2025Mambavision), such as Mambavision-S-1K and Mambavision-L-21K in half-precision (16-bit floating-point or FP16) format trained on ImageNet data (deng2009imagenet) to showcase multi-modal effectiveness of RAMBO. Furthermore, we assess RAMBO on Hymba (dong2024hymba), a hybrid-head parallel architecture that combines transformer attention mechanisms and SSMs, to show the versatility of the proposed framework. We report both perplexity (on WikiText (hu2024can)) and accuracy as evaluation metrics. Perplexity, defined as the exponential of the average negative log-likelihood over a sequence, measures predictive capability (hu2024can), whereas accuracy quantifies the proportion of correct predictions.
Table 1. RAMBO evaluation on various models and datasets.
@@ 5.3. Results
This segment presents the degradation of model performance across benchmarks achieved by RAMBO (Table 1). The attributes of each model, specifically the model name and parameter count, are presented in the firts column. The second column presents the count of bit-flips injected to induce performance degradation. Subsequent columns furnish benchmark results, before and after the attack, for each task. All models demonstrated strong baseline performance, with LAMBADA accuracies up to 82.20%, and WikiText perplexities between 16.31 and 26.69. Similarly, high accuracies were observed across reasoning and commonsense benchmarks such as ARC-Easy (53.12– 76.94%), HellaSwag (43.04–62.69%), PIQA (67.03–77.31%), and Winogrande (51.46– 66.61%). “NA” indicates when the benchmark is incompatible and does not apply to the model.
Following the injection of only a few targeted bit-flips, RAMBO produced severe accuracy degradation across all tasks. In the Mamba-1.4b model, a single bit-flip caused complete collapse on LAMBADA (accuracy dropped from 74.64% to 0%) and raised perplexity from 18.94 to 3.75 × 106, while also reducing ARC-Easy accuracy from 62.5% to 14%. The Mamba-2.8b model exhibited similar vulnerability, where one bit-flip reduced LAMBADA accuracy from 78.34% to 4% and increased perplexity from 16.31 to 1.23 × 105. Even smaller models, such as Mamba-370M, experienced notable drops, with four bit-flips lowering Arc-Easy accuracy from 53.125% to 28.67% and increasing perplexity from 24.87 to 1.26 $\times 10^{11}$.
@@ 5.4. Attack Efficacy across Quantization Levels
Our evaluation has, up to this point, concentrated on RAMBO using half-precision FP16 models. However, a model’s vulnerability to bit-flip attacks is largely determined by quantization precision. Since diverse numerical formats display varying sensitivities to perturbations, the effect of a bit-flip, along with the quantity of flips needed to cause model disruption, can differ significantly across different quantizations (das2024attentionbreaker). Furthermore, previous BFA defenses commonly suggest quantization as a strategy for mitigation, necessitating the evaluation of RAMBO under quantized conditions (rakin2021t).
To this end, we evaluate RAMBO on INT4 (w4a16) and INT8 (w8a8) variants of Mamba-1.4b and Mamba-2.8b. Our findings show that these quantized models remain highly vulnerable to bit-flip attacks. Notably, a single bit-flip is sufficient to reduce the accuracy of the Mamba-2.8b w4a16 model on PIQA from 73.29% to 25.33% (refer Table 1), demonstrating that RAMBO remains extremely effective even under aggressive quantization.
@@ 5.5. Attack Efficacy across Task Modality
To assess RAMBO’s effectiveness beyond NLP settings, we applied it to FP16 Mamba-Vision models with 50M–200M parameters trained on ImageNet. Our results show that RAMBO remains highly effective even in vision tasks, successfully identifying critical bit positions whose perturbation leads to substantial performance degradation. For example, in the MambaVision-L-21K model, RAMBO identified 24 critical bits that, when flipped, reduced the model’s accuracy from 86.1% to 47.2% (refer Table 1). This demonstrates that RAMBO generalizes effectively beyond NLP and remains a potent bit-flip attack methodology for Mamba-based architectures on vision tasks.
@@ 5.7. Gradient-free Attack Results
We evaluate RAMBO in a fully gradient-free setting, where gradient information is unavailable throughout the attack. Parameter sensisitivity scores are derived solely from weight magnitudes by setting $\alpha=0$ in Equation 7. Under this configuration, only a single bit-flip on the 8-bit Quamba-2.8B-W8A8 model reduces ARC-Easy accuracy from 65.26% to 0%, demonstrating that magnitude-based scoring remains highly effective in identifying critical bits. This flexibility allows adversaries to select either gradient-free or gradient-based modes depending on their goals and computational constraints.
Notably, many existing BFA defenses rely on restricting or obfuscating gradient access. RAMBO circumvents such defenses entirely by operating on magnitude-based importance alone. This adaptability underscores RAMBO’s robustness in gradient-restricted environments and establishes it as a resilient BFA framework.
@@ 5.8. Gray-box Attack Results
To evaluate RAMBO’s effectiveness under grey-box scenarios with partial access to model parameters, we simulate such a condition by restricting access to a only a subset of model layers, specifically the final two Mamba blocks in the Mamba-1.4b model. Even with this limited access, RAMBO remains highly potent. On the Arc-Easy dataset, flipping only a single bit in this partially visible Mamba-1.4b model is sufficient to reduce accuracy from 62.5% to 15.2%, demonstrating the strength of our approach under constrained access.
@@ 5.9. Comparison with State-of-the-art LLM Bit-flip Attacks
Although RAMBO is the first SSM-aware BFA and thus lacks a direct SSM-specific baseline, we assess how its exclusionary optimization strategy compares with LLM-based BFAs. To this end, we evaluate it against GenBFA optimization employed in the AttentionBreaker framework (das2024attentionbreaker). As shown in Figure 4(b), we observe that both methods identify the same 22 critical bits in Mambavision-S-1K, but RAMBO converges significantly faster, within 26 iterations (green line) compared to GenBFA’s 70 (blue line), demonstrating superior efficiency despite its simpler design.
Prior works, such as AttentionBreaker (das2024attentionbreaker) and SBFA (guo2025sbfa), show that single or few bit-flips can collapse models with billions of parameters. Similarly, RAMBO achieves catastrophic degradation with just one bit-flip in Mamba-1.4b (FP16) and Quamba-2.8b (INT4). Even the hybrid Hymba architecture is oberserved to be extremely vulnerable and its performance colapses with merely two bit-flips. These results highlight the severe security risk posed by bit-flip attacks in both transformer- and state-space-based models.
@@ 6. Conclusion
This research investigates the vulnerability of state-space based models like Mamba and presents RAMBO, a novel BFA framework targeting these models. RAMBO introduces a novel sensitivity analysis of state space model architectures, thereby resulting in compromised performance. Specifically, perturbing merely one-bit (7.14 x 10-10% of the model parameters) in Mamba-1.4b model results on LAMBADA dataset prediction scores dropping from 74.64% to 0%. The results highlight RAMBO’s utility in demonstrating Mamba’s vulnerability to such adversarial interventions.

### v2

# COBRA: Catastrophic Bit-flip Reliability Analysis of State-Space Models

@@ Abstract.
Among such threats, hardware-induced bit-flip attacks (BFAs) pose a particularly severe risk by corrupting model parameters through memory faults, thereby undermining model accuracy and functional integrity. To investigate this vulnerability, we introduce COBRA, the first BFA framework specifically designed to target Mamba-based architectures. Through experiments on the Mamba-1.4b model with LAMBADA benchmark—a cloze-style word- prediction task—we demonstrate that flipping merely a single critical bit can catastrophically reduce accuracy from 74.64% to 0% and increase perplexity from 18.94 to 3.75 × 106. These results demonstrate the pronounced fragility of SSMs to adversarial perturbations. The framework is open-sourced at https://anonymous.4open.science/r/RAMBO-DA22.
@@ 1. Introduction
State-space architectures such as Mamba implement fundamentally different information-propagation and parameterization mechanisms, trade off recurrence and selectivity for attention mechanisms (gu2023Mamba). Due to these architectural distinctions, it is not appropriate to assume that attack strategies and defenses developed for transformers transfer directly to SSMs. Consequently, the absence of a systematic study of BFAs on SSMs constitutes a substantive gap in the current AI hardware robustness and security literature. To address this gap, we propose the framework “COBRA” — a first of its kind SSM-aware BFA pipeline. The primary contribution of the paper are as follows:
  • We identify and formalize a previously unexplored vulnerability of state-space models (e.g. Mamba) to bit-flip attacks, and propose COBRA, a first of its kind SSM-aware attack approach, bridging the gap between BFA research and structured sequence models.
  • COBRA leverages the structural properties of Mamba-style SSM layers to prioritize critical parameter regions, adapts gradient-estimation and search heuristics, and uses the resulting perturbation effects to identify minimal bit-flip sets that maximally disrupt model behavior.
  • COBRA uncovers a significant vulnerability of Mamba models. A mere one bit-flip (7.14 × 10-10% of all bits) in Mamba-1.4b, can reduce the LAMBADA word prediction accuracy from 74.64% to 0%, while increasing Wikitext perplexity from 18.94 to 3.75 × 106.
@@ 3. Threat Model
In this work, we consider both white-box and gray-box threat models. The white-box setting applies because the targeted models are open-sourced, granting full access to their architecture and parameters. The gray-box setting reflects scenarios in which the adversary has only partial visibility into model internals. Within this setting, we focus on untargeted attacks, which aim to induce graceless degradation in overall model performance rather than targeting specific outputs. Such attacks are particularly insidious, as they compromise accuracy across a broad range of inputs while avoiding distinct failure patterns, making them more challenging to detect and defend against than targeted attacks (codematch; qian2023survey; nazari2024forget). Therefore, COBRA strategically employs an untargeted attack to maximize disruption and degrade model performance. This attack also fundamentally differs from denial of service attacks, which aim to overwhelm system resources to render services unavailable.
4. Proposed COBRA Methodology
@@ 4.1. Theoretical Vulnerability Analysis
Here, we present parameter-ranking procedure to identify plausible targets for the COBRA attack. We rank the SSM parameters’ based on their importance as follows:
@@ 5.1. Experimental Setup
We evaluated COBRA on a diverse set of models, including Mamba and Mamba2, ranging from 370 million and 2.8 billion parameters (Mamba; Mamba2). We further assess COBRA on quantized models such as Quamba-1.4b-w8a8 (8-bit weights and 8-bit activations), Quamba-1.4b-w4a16, Quamba-2.8B-w8a8, and Quamba-2.8b-w4a16 to evaluate its effectiveness under low-precision settings. Model performance was assessed using standard benchmarks, such as the tasks from the Language Model Evaluation Harness (eval-harness), including Arc-Easy, HellaSwag, PIQA and Winograde, which probe reasoning and generalization across a variety of domains. Additionally, we tested on the LAMBADA dataset (paperno2016lambada), a word-prediction/ cloze-style natural language understanding task. We extended our evaluation to include Vision-Mamba models (hatamizadeh2025Mambavision), such as Mambavision-S-1K and Mambavision-L-21K in half-precision (16-bit floating-point or FP16) format trained on ImageNet data (deng2009imagenet) to showcase multi-modal effectiveness of COBRA. Furthermore, we assess COBRA on Hymba (dong2024hymba), a hybrid-head parallel architecture that combines transformer attention mechanisms and SSMs, to show the versatility of the proposed framework. We report both perplexity (on WikiText (hu2024can)) and accuracy as evaluation metrics. Perplexity, defined as the exponential of the average negative log-likelihood over a sequence, measures predictive capability (hu2024can), whereas accuracy quantifies the proportion of correct predictions.
Table 1. COBRA evaluation on various models and datasets.
@@ 5.3. Results
This segment presents the degradation of model performance across benchmarks achieved by COBRA (Table 1). The attributes of each model, specifically the model name and parameter count, are presented in the firts column. The second column presents the count of bit-flips injected to induce performance degradation. Subsequent columns furnish benchmark results, before and after the attack, for each task. All models demonstrated strong baseline performance, with LAMBADA accuracies up to 82.20%, and WikiText perplexities between 16.31 and 26.69. Similarly, high accuracies were observed across reasoning and commonsense benchmarks such as ARC-Easy (53.12– 76.94%), HellaSwag (43.04–62.69%), PIQA (67.03–77.31%), and Winogrande (51.46– 66.61%). “NA” indicates when the benchmark is incompatible and does not apply to the model.
Following the injection of only a few targeted bit-flips, COBRA produced severe accuracy degradation across all tasks. In the Mamba-1.4b model, a single bit-flip caused complete collapse on LAMBADA (accuracy dropped from 74.64% to 0%) and raised perplexity from 18.94 to 3.75 × 106, while also reducing ARC-Easy accuracy from 62.5% to 14%. The Mamba-2.8b model exhibited similar vulnerability, where one bit-flip reduced LAMBADA accuracy from 78.34% to 4% and increased perplexity from 16.31 to 1.23 × 105. Even smaller models, such as Mamba-370M, experienced notable drops, with four bit-flips lowering Arc-Easy accuracy from 53.125% to 28.67% and increasing perplexity from 24.87 to 1.26 $\times 10^{11}$.
@@ 5.4. Attack Efficacy across Quantization Levels
Our evaluation has, up to this point, concentrated on COBRA using half-precision FP16 models. However, a model’s vulnerability to bit-flip attacks is largely determined by quantization precision. Since diverse numerical formats display varying sensitivities to perturbations, the effect of a bit-flip, along with the quantity of flips needed to cause model disruption, can differ significantly across different quantizations (das2024attentionbreaker). Furthermore, previous BFA defenses commonly suggest quantization as a strategy for mitigation, necessitating the evaluation of COBRA under quantized conditions (rakin2021t).
To this end, we evaluate COBRA on INT4 (w4a16) and INT8 (w8a8) variants of Mamba-1.4b and Mamba-2.8b. Our findings show that these quantized models remain highly vulnerable to bit-flip attacks. Notably, a single bit-flip is sufficient to reduce the accuracy of the Mamba-2.8b w4a16 model on PIQA from 73.29% to 25.33% (refer Table 1), demonstrating that COBRA remains extremely effective even under aggressive quantization.
@@ 5.5. Attack Efficacy across Task Modality
To assess COBRA’s effectiveness beyond NLP settings, we applied it to FP16 Mamba-Vision models with 50M–200M parameters trained on ImageNet. Our results show that COBRA remains highly effective even in vision tasks, successfully identifying critical bit positions whose perturbation leads to substantial performance degradation. For example, in the MambaVision-L-21K model, COBRA identified 24 critical bits that, when flipped, reduced the model’s accuracy from 86.1% to 47.2% (refer Table 1). This demonstrates that COBRA generalizes effectively beyond NLP and remains a potent bit-flip attack methodology for Mamba-based architectures on vision tasks.
@@ 5.7. Gradient-free Attack Results
We evaluate COBRA in a fully gradient-free setting, where gradient information is unavailable throughout the attack. Parameter sensisitivity scores are derived solely from weight magnitudes by setting $\alpha=0$ in Equation 7. Under this configuration, only a single bit-flip on the 8-bit Quamba-2.8B-W8A8 model reduces ARC-Easy accuracy from 65.26% to 0%, demonstrating that magnitude-based scoring remains highly effective in identifying critical bits. This flexibility allows adversaries to select either gradient-free or gradient-based modes depending on their goals and computational constraints.
Notably, many existing BFA defenses rely on restricting or obfuscating gradient access. COBRA circumvents such defenses entirely by operating on magnitude-based importance alone. This adaptability underscores COBRA’s robustness in gradient-restricted environments and establishes it as a resilient BFA framework.
@@ 5.8. Gray-box Attack Results
To evaluate COBRA’s effectiveness under grey-box scenarios with partial access to model parameters, we simulate such a condition by restricting access to a only a subset of model layers, specifically the final two Mamba blocks in the Mamba-1.4b model. Even with this limited access, COBRA remains highly potent. On the Arc-Easy dataset, flipping only a single bit in this partially visible Mamba-1.4b model is sufficient to reduce accuracy from 62.5% to 15.2%, demonstrating the strength of our approach under constrained access.
@@ 5.9. Comparison with State-of-the-art LLM Bit-flip Attacks
Although COBRA is the first SSM-aware BFA and thus lacks a direct SSM-specific baseline, we assess how its exclusionary optimization strategy compares with LLM-based BFAs. To this end, we evaluate it against GenBFA optimization employed in the AttentionBreaker framework (das2024attentionbreaker). As shown in Figure 4(b), we observe that both methods identify the same 22 critical bits in Mambavision-S-1K, but COBRA converges significantly faster, within 26 iterations (green line) compared to GenBFA’s 70 (blue line), demonstrating superior efficiency despite its simpler design.
Prior works, such as AttentionBreaker (das2024attentionbreaker) and SBFA (guo2025sbfa), show that single or few bit-flips can collapse models with billions of parameters. Similarly, COBRA achieves catastrophic degradation with just one bit-flip in Mamba-1.4b (FP16) and Quamba-2.8b (INT4). Even the hybrid Hymba architecture is oberserved to be extremely vulnerable and its performance colapses with merely two bit-flips. These results highlight the severe security risk posed by bit-flip attacks in both transformer- and state-space-based models.
@@ 6. Conclusion
This research investigates the vulnerability of state-space based models like Mamba and presents COBRA, a novel BFA framework targeting these models. COBRA introduces a novel sensitivity analysis of state space model architectures, thereby resulting in compromised performance. Specifically, perturbing merely one-bit (7.14 x 10-10% of the model parameters) in Mamba-1.4b model results on LAMBADA dataset prediction scores dropping from 74.64% to 0%. The results highlight COBRA’s utility in demonstrating Mamba’s vulnerability to such adversarial interventions.